# Free field construction of Heterotic string compactified on Calabi-Yau manifolds of Berglund-Hubsch type in the Batyrev-Borisov combinatorial approach

### v5

@@ 1 Introduction
We first present Borisov’s construction for $N=2$ SCFT models corresponding to Calabi-Yau manifolds of Berglund-Hubsch type. Then, in the left moving sector of the model we find the generators of $N=1$ spacetime supersymmetry and choose the set of left-movers that are mutually local with them. After this, in the right sector we find the additional generators of $E(6)$ symmetry and choose the set of right movers that are mutually local with them. At the last step, we find the products of the left- and right-moving vertices of the states obtained in this way, which are mutually local to each other.
@@ 2 The original Conformal field theory
$T^{L}(z)=\frac{1}{2}\eta_{\mu\nu}\partial x^{\mu}\partial x^{\nu}+\frac{1}{2}\eta_{\mu\nu}\psi^{\mu}(z)\partial\psi^{\nu}(z)+T^{L}_{CY}(z).$ (1)
@@
$T^{R}(\bar{z})=\frac{1}{2}\eta_{\mu\nu}\bar{\partial}\bar{x}^{\mu}\bar{\partial}\bar{x}^{\nu}+\frac{1}{2}(\bar{\partial}Y_{I})^{2}+\frac{1}{2}(\bar{\partial}\Phi_{\alpha})^{2}+T^{R}_{CY}(\bar{z}).$ (2)
@@ 4 Berglund-Hubsch type CY manifolds and N=2 SCFT models
Following BOR1 ; BOR2 we define the Vertex algebra of Heterotic string as the cohomology with respect to the set diffentials $D_{\vec{m}}$ and $D_{\vec{n}}$
@@
For shortness, we will call the set of states belonging to the intersection of the set of BRST cohomology defined by the differential $Q_{BRST}$ and the set of Borisov cohomology BOR1 ; BOR2 defined by the differentials $D_{m}$ and $D_{n}$, the space of quasi-physical states.
@@ 6 Massless left movers and N=1 Space-Time supersymmetry
The vertices of massless physical states are interesting because they play two roles.The first is that it is the states of this set that must correspond to the observed elementary particles. Secondly, some of these vertices can be used to extend the symmetry of the theory. They can be taken as currents whose integrals become additional generators for this extension.
By imposing the requirements on the vertices to be simultaneously $Q_{BRST}$, $D_{\vec{m}}$ and $D_{\vec{n}}$ cohomology, we find the following massless left vertices in the NS-subsector with canonical picture number (1)(-1) and in the R-subsector sector with canonical picture number $(-\frac{1}{2})$.
In the NS-subsector we get the left-moving vertex of massless vector boson
@@
In the NS-subsector we also find the left-moving vertex of massless scalar bosons
@@
where $\vec{m}\in\Delta^{+}$ and $\vec{n}\in\Delta^{-}$. However, these vertices are exact with respect to Borisov differentials (14) and thus they do not participate in the construction of the physical states.
@@
The first two currents $J^{+}_{\sigma}$ and $J^{-}_{\dot{\sigma}}$ are mutually local, as are the other two currents. We will use the first pair to extend Poincaré symmetry to $N=1$ spacetime supersymmetry.
@@
The supercharges $\cal{Q}_{\sigma}$ and ${\cal{Q}}_{\dot{\sigma}}$, which are spinors with respect to the Poincaré algebra, form together with the generators of this algebra, $P_{\mu}$ and $J_{\mu\nu}$, the $N=1$ Poincaré superalgebra .
In order to obtain $N=1$ spacetime supersymmetry in the theory, we must leave from the vertices of $V^{\vec{L}}_{\vec{\mu}_{L}}$, where $\vec{\mu}_{L}=(q,\ \vec{\lambda},\ Q^{L}_{CY})$, which are the cohomologies of $Q_{BRST}$, $D_{m}$, $D_{n}$ only those vertices that are mutually local with $J^{+}_{\sigma}$ and $J^{-}_{\dot{\sigma}}$.
@@
From the GSO equation it follows that the total internal charges $Q^{L}_{CY}$ of the vertices are integers or half-integers.
This can be obtained more explicitly if we consider that the vectors $\vec{\lambda}$ are weights of the algebra $SO(1,3)$, belonging to one of the four conjugacy classes of the weight lattice.
@@
From GSO equation it follows that in NS sector the vectors $\vec{\lambda}$ fall into classes $[0]$ and $[V]$ and in R sector the vector $\vec{\lambda}$ fall into classes[S][S] and $[C]$ of $SO(1,3)$.
As a result, we obtain the following connection between the conjugacy class of $\vec{\lambda}$ and the $U(1)$ charge of the CY factor for all 4 cases of left vertices.
@@ 7 Right-moving sector, N=0 CFT
In order to get critical string in the right-moving sector we should add the bosonic fields $Y_{I}(\bar{z})$, I=1,,8I=1,...,8 which are compactified on the torus of the algebra $E(8)$ with the central charge $8$ and the bosons $\Phi_{\alpha}(\bar{z})$, compactified on the torus of the algebra $SO(10)$ with the central charge $5$. The final contribution to the critical dimension is given by the right-moving part of the compact Calabi-Yau subsector with the central charge $9$.
The right-moving energy-momentum tensor is
@@
where $\vec{\eta}$ is a vector of the $E(8)$ root lattice, and $\vec{L}$ is the vector of the $SO(10)$ weight lattice.
Among the vertices (55) we find, as in the left sector, BRST-invariant massless ones. First of all this is $SO(1,3)$ vector $V^{\mu}(\bar{z})=\imath\bar{\partial}\bar{X}^{\mu}(\bar{z})$. Then we also find the currents of $E(8)$ algebra
$J^{I}(\bar{z})=\imath\bar{\partial}\bar{Y}^{I}(\bar{z}),\ I=1,...,8,\ J_{\vec{\epsilon}}(\bar{z})=\exp{[\imath\epsilon_{I}\bar{Y}^{I}]}(\bar{z}),\ \vec{\epsilon}^{2}=2,$ (43)
@@
Also there are the currents of $SO(10)$ algebra:
where the vectors $\vec{\rho}$
$\vec{\rho}=(\pm 1,\pm 1,0,0,0)+\text{permutations}$ (46)
are the roots of $SO(10)$. And the current of $U(1)$ algebra of the right Calabi-Yau sector
@@
In the right-moving sector there are also massless $SO(10)$ spinors
$J^{\pm}_{\omega}(\bar{z})=\exp{(\imath\omega_{\alpha}\Phi_{\alpha}\pm\frac{1}{2}\bar{H}_{CY})(\bar{z})}$ (48)
$J^{\pm}_{\dot{\omega}}(\bar{z})=\exp{(\imath\dot{\omega}_{\alpha}\Phi_{\alpha}\pm\frac{1}{2}\bar{H}_{CY})(\bar{z})},$
The currents $J^{+}_{\omega}$ is mutually local with $J^{-}_{\dot{\omega}}$, and $J^{-}_{\omega}$ is mutually local with $J^{+}_{\dot{\omega}}$.
We also emphasize that, as it can be easily verified, all these currents are cohomologies of the differentials $D_{\vec{m}}$ and $D_{\vec{n}}$. By choosing one of these pairs (we choose $J^{+}_{\omega}(\bar{z})$, $J^{-}_{\dot{\omega}}$), we can extend the product of the algebra $SO(10)\times U(1)$ to the algebra $E(6)$, adding them to the currents of the algebra $SO(10)$ and the algebra $U(1)$.
One can rewrite the $E_{6}$ currents in terms of simple roots of $E_{6}$
$\vec{\alpha}_{i}=e_{1}-e_{2},\ .,\ \vec{\alpha}_{4}=e_{4}-e_{5},\ \vec{\alpha}_{5}=e_{4}+e_{5},$ (49)
$\vec{\alpha}_{6}=-\frac{1}{2}(e_{1}+.+e_{5})+\frac{\sqrt{3}}{2}e_{6},$
where $e_{i}$ are the orthonormal basic vectors in $\mathbb{R}^{6}$. So that the Cartan subalgebra currents are
$J_{j}(\bar{z})=\imath\vec{\alpha}_{j}\cdot\bar{\partial}\vec{\bar{H}}(\bar{z}),\ j=1,.,6,\text{where}$ (50)
$\vec{\bar{H}}(\bar{z})=(\bar{H}_{1}(\bar{z}),.,\bar{H}_{5}(\bar{z}),\bar{H}_{6}(\bar{z})),$
[…]
The currents of ladder $E(6)$ operators are given by
$E_{j}(\bar{z})=\exp{[\imath\vec{\alpha}_{j}\vec{\bar{H}}]}(\bar{z}),$ (51)
$F_{j}(\bar{z})=\exp{[-\imath\vec{\alpha}_{j}\vec{\bar{H}}]}(\bar{z}),\ j=1,.,6.$
The other $E(6)$ currents can be generated from (50), (51) by the OPE’s.
$E(6)$ is known to be considered as a possible Grand Unified gauge group, which after spontaneous breaking gives rise to the $SU(3)\times SU(2)\times U(1)$ gauge group of the Standard Model.
It should be noted that it is the appearance of the $E(6)$ symmetry in the right sector that ensures the self-consistency of the theory after the appearance of $N=1$ SUSY in the left sector, which (i.e. $N=1$ SUSY) is also considered a necessary element of the Grand Unified Theory.
In the right sector we also find a massless boson vertex that belongs to the one-dimensional representation of $SO(10)$
$V^{[1]}_{\vec{m}}=\exp{(\bar{H}_{CY}+\vec{m}\cdot\vec{X}^{-})},$ (52)
another massless boson vertex that belongs to the 10-dimensional representation of $SO(10)$
$V^{[10]}_{\vec{m}}=\exp{(i\vec{\Lambda}\cdot\vec{\Phi}+\vec{m}\cdot\vec{X}^{-})},$ (53)
and one more massless boson vertex that belongs to the 16-dimensional spinor representation of $SO(10)$
$V^{[16]}_{\vec{m}}=\exp{(i(\vec{\Lambda}+\vec{\omega})\cdot\vec{\Phi}+\frac{1}{2}\bar{H}_{CY}+\vec{m}\cdot\vec{X}^{-})},$ (54)
where $\vec{m}\in\Delta^{+}$, $\vec{\Lambda}=(\pm 1,0,0,0,0)+\text{permutations}$ and $\vec{\Lambda}+\vec{\omega}=\vec{\dot{\omega}}$ .
These three types of vertices are the cohomologies of the differentials $D_{\vec{m}}$ and $D_{\vec{n}}$ and, as will be shown below, form a 27-dimensional representation of E(6).
The right massless sector also contains $\bar{27}$ representations of E(6), which are similarly constructed from dual representations of $SO(10)$ whose vertices depend on $\vec{n}\in\Delta^{-}$. The numbers of representations $27$ and $\bar{27}$ are equal to the number of points in the reflexive Batyrev polytopes for a given CY-manifold.
Finally, the fourth class of massless right vertices, which are also BRST, $D_{\vec{m}}$ and $D_{\vec{n}}$ cohomology, are $E(8)\times E(6)$ singlets
@@
$\Delta(\vec{S},\vec{m},\vec{n})=\frac{\vec{S}^{2}}{2}+\vec{m}\cdot\vec{n}+\frac{\vec{m}\cdot\vec{a}^{-}+\vec{}\cdot\vec{a}^{-}}{2}=1,$ (56)
@@
There are three types of solutions to all the above requirements that are imposed on the $E(8)\times E(6)$ singlets.
The first type of the singlet vertices is
$V(\vec{S},\vec{m},\vec{n})=\ \exp{(\vec{m}\cdot\vec{X}^{-}+\vec{n}\cdot\vec{X}^{+}+\imath p_{\mu}\bar{X}^{\mu})(\bar{z})}$ (58)
, where the vectors $\vec{m}\in\Delta^{+}$, $\vec{n}\in\Delta^{-}$ and their product $\vec{m}\cdot\vec{n}$ is equal to 0.
The second type of the singlet vertices is
$V(\vec{S},\vec{m},\vec{n})=\ \sum_{i}m_{i}\exp{(-i\bar{H}_{i}+\vec{m}\cdot\vec{X}^{-}+\imath p_{\mu}\bar{X}^{\mu})(\bar{z}))},$ (59)
@@
The third type of the singlet vertices is
$V(\vec{S},\vec{m},\vec{n})=\ \sum_{i}n_{i}\exp{(i\bar{H}_{i}+\vec{n}\cdot\vec{X}^{+}+\imath p_{\mu}\bar{X}^{\mu})(\bar{z}))}$ (60)
@@
The total number of singlet massless states determined by these conditions in the case of Quintic coincides with the result of D. Gepner Gep .

### v6

@@ 1 Introduction
We first present Borisov’s construction for $N=2$ SCFT models corresponding to Calabi-Yau manifolds of Berglund-Hubsch type. Then, in the left moving sector of the model we find the generators of $N=1$ spacetime supersymmetry and choose the set of left-movers that are mutually local with them. After this, in the right sector we find the additional generators of $E(6)$ symmetry and choose the set of right movers that are mutually local with them.
At the last step, we find the products of the left- and right-moving vertices of the states obtained in this way, which are mutually local to each other and permit to achieve modular invariance in constructing heterotic string theory compactified on CY-manifolds of BH type.
@@ 2 The original Conformal field theory
$T^{L}(z)=-\frac{1}{2}\eta_{\mu\nu}\partial x^{\mu}\partial x^{\nu}-\frac{1}{2}\eta_{\mu\nu}\psi^{\mu}(z)\partial\psi^{\nu}(z)+T^{L}_{CY}(z).$ (1)
@@
$T^{R}(\bar{z})=-\frac{1}{2}\eta_{\mu\nu}\bar{\partial}\bar{x}^{\mu}\bar{\partial}\bar{x}^{\nu}+\frac{1}{2}(\bar{\partial}Y_{I})^{2}+\frac{1}{2}(\bar{\partial}\Phi_{\alpha})^{2}+T^{R}_{CY}(\bar{z}).$ (2)
@@ 4 Berglund-Hubsch type CY manifolds and N=2 SCFT models
Following to Borisov BOR1 ; BOR2 we define the Vertex algebra of Heterotic string as the cohomology with respect to the diffential $D_{\vec{1},\vec{1}}$ defined as the sum of the differntials $D_{\vec{m}}$ and $D_{\vec{n}}$, where
@@
For shortness, we will call the set of states belonging to the intersection of the set of BRST cohomology defined by the differential $Q_{BRST}$ and the set of Borisov cohomology BOR1 ; BOR2 defined by the differential $D_{\vec{1},\vec{1}}$, the space of quasi-physical states.
@@ 6 Massless left movers and N=1 Space-Time supersymmetry
The vertices of massless physical states are interesting because they play two roles.The first is that it is the states of this set that must correspond to the observed elementary particles.
Secondly, some of these vertices can be used to extend the symmetry of the theory. They can be taken as currents whose integrals become additional generators for this extension.
By imposing the requirements on the vertices to be simultaneously $Q_{BRST}$ and $D_{\vec{1},\vec{1}}$ cohomology, we find the following massless left vertices in the NS-subsector with canonical picture number (1)(-1) and in the R-subsector sector with canonical picture number $(-\frac{1}{2})$.
In the NS subsector we obtain a left-hand vertex belonging to the $N=1$ Super-Poincaré vector representation of the space-time symmetry
@@
In the NS-subsector we also find a left-hand vertex belonging to the $N=1$ Super-Poincaré singlet representation of the space-time symmetry
@@
where $\vec{m}\in\Delta^{+}$ and $\vec{n}\in\Delta^{-}$.
@@
The first two currents $J^{+}_{\sigma}$ and $J^{-}_{\dot{\sigma}}$ are mutually local, as are the other two currents. We will use the first pair to extend Poincaré symmetry to $N=1$ space-time supersymmetry.
@@
The supercharges $\cal{Q}_{\sigma}$ and ${\cal{Q}}_{\dot{\sigma}}$, which are spinors with respect to the Super-Poincaré algebra, form together with the generators of this algebra, $P_{\mu}$ and $J_{\mu\nu}$, the $N=1$ Poincaré superalgebra .
In order to obtain $N=1$ space-time supersymmetry in the theory, we must leave from the vertices of $V^{\vec{L}}_{\vec{\mu}_{L}}$, where $\vec{\mu}_{L}=(q,\ \vec{\lambda},\ Q^{L}_{CY})$, which are the cohomologies of $Q_{BRST}$ and $D_{\vec{1},\vec{1}}$, only those vertices that are mutually local with $J^{+}_{\sigma}$ and $J^{-}_{\dot{\sigma}}$.
@@
From the GSO equation it follows that the total internal charges $Q^{L}_{CY}$ of the vertices are integers or half-integers. This can be obtained more explicitly if we consider that the vectors $\vec{\lambda}$ are weights of the algebra $SO(1,3)$, belonging to one of the four conjugacy classes of the weight lattice.
@@
From GSO equation it follows that in NS sector the vectors $\vec{\lambda}$ fall into classes $[0]$ and $[V]$ and in R sector the vector $\vec{\lambda}$ fall into classes[S][S] and $[C]$ of $SO(1,3)$. As a result, we obtain the following connection between the conjugacy class of $\vec{\lambda}$ and the $U(1)$ charge of the CY factor for all 4 cases of left vertices.
@@ 7 Right-moving sector, N=0 CFT
In order to obtain a critical string with the total central charge of $26$ in the right-handed sector, we add the boson fields $Y_{I}(\bar{z})$, I=1,,8I=1,...,8, compactified onto the torus of the algebra $E(8)$ with a central charge of $8$ and the bosons $\Phi_{\alpha}(\bar{z})$, compactified onto the torus of the algebra $SO(10)$ with a central charge of $5$. The final contribution to the critical dimension is given by the right-moving part of the compact Calabi-Yau subsector with the central charge $9$.
The full right-moving energy-momentum tensor is
@@
where $\vec{\eta}$ is a vector of the $E(8)$ root lattice, and $\vec{\Lambda}$ is the vector of the $SO(10)$ weight lattice.
Similar to what was done in the left sector, we impose the requirements that the right moving vertices be both $BRST$ and $D_{\vec{1},\vec{1}}$ cohomologies, and also respect symmetries under $E(8)$ and $E(6)$ symmetries.
As in the left-handed sector, the right-handed vertices of massless physical states are interesting because they play two roles. First, it is the states of this set that must correspond to observable elementary particles. Second, some of these vertices can be used to extend the symmetry of the theory. As a result, it turns out that, as in the left-hand sector, we find four types of massless vertices that satisfy our requirements.
The first type $[0]$ consists of the following $SO(10)$ singlets (and of their $E(6)$ partners).
It includes $SO(1,3)$ vector $\imath\bar{\partial}\bar{X}^{\mu}(\bar{z})$ and the currents of $SO(10)$
where the vectors $\vec{\rho}$ are the roots of $SO(10)$
$\vec{\rho}=(\pm 1,\pm 1,0,0,0)+\text{permutations};$ (44)
It also includes the currents of $E(8)$ algebra
$J^{I}(\bar{z})=\imath\bar{\partial}\bar{Y}^{I}(\bar{z}),\ I=1,...,8,\ J_{\vec{\epsilon}}(\bar{z})=\exp{[\imath\epsilon_{I}\bar{Y}^{I}]}(\bar{z}),\ (\vec{\epsilon})^{2}=2,$ (45)
@@
The current of $U(1)$ algebra of the right Calabi-Yau sector
@@
also belong to $[0]$.
Also, there is another set of $[0]$: the massless right vertices, which also consist of BRST and Borisov cohomology and are $E(8)\times E(6)$ singlets depending on the Calabi-Yau data.
@@
$\Delta(\vec{S},\vec{m},\vec{n})=\frac{\vec{S}^{2}}{2}+\vec{m}\cdot\vec{n}+\frac{\vec{m}\cdot\vec{a}^{-}+\vec{n}\cdot\vec{a}^{-}}{2}=1,$ (49)
@@
There are three cases of solutions to all the above requirements that are imposed on the $E(8)\times E(6)$ singlets.
The first case is
$V(\vec{S},\vec{m},\vec{n})=\ \exp{(\vec{m}\cdot\vec{X}^{-}+\vec{n}\cdot\vec{X}^{+}+\imath p_{\mu}\bar{X}^{\mu})(\bar{z})};$ (51)
where the vectors $\vec{m}\in\Delta^{+}$, $\vec{n}\in\Delta^{-}$ and their product $\vec{m}\cdot\vec{n}$ is equal to 0.
The second case of the singlet of $SO(10)$ vertices is
$V(\vec{S},\vec{m},\vec{n})=\ \sum_{i}m_{i}\exp{(-i\bar{H}_{i}+\vec{m}\cdot\vec{X}^{-}+\imath p_{\mu}\bar{X}^{\mu})(\bar{z}))};$ (52)
@@
The third case is
$V(\vec{S},\vec{m},\vec{n})=\ \sum_{i}n_{i}\exp{(i\bar{H}_{i}+\vec{n}\cdot\vec{X}^{+}+\imath p_{\mu}\bar{X}^{\mu})(\bar{z}))}.$ (53)
@@
The total number of singlet massless states depends on and is determined by the CY data. In the case of Quintic their number coincides with the result of D. Gepner Gep .
The second type $[V]$ of the massless vertices consists of the following vertices of $10$-dimensional vector representation of $SO(10)$ (and their $E(6)$ partner).
$V^{[10]}_{\vec{m}}=\exp{(i\vec{\Lambda}\cdot\vec{\Phi}+\vec{m}\cdot\vec{X}^{-})}.$ (54)
The third type $[S]$ consists of the following vertices massless $SO(10)$ spinors (and their $E(6)$ partners)
$J^{+}_{\omega}(\bar{z})=\exp{(\imath\omega_{\alpha}\Phi_{\alpha}+\frac{1}{2}\bar{H}_{CY})(\bar{z})}$ (55)
$\omega_{\alpha}=\pm\frac{1}{2},\ \sum\omega_{\alpha}=\frac{1}{2}\mod 2\mathbb{Z}.$
The fourth type $[C]$ consists of the following vertices massless $SO(10)$ spinors (and their $E(6)$ partners).
$J^{-}_{\dot{\omega}}(\bar{z})=\exp{(\imath\dot{\omega}_{\alpha}\Phi_{\alpha}-\frac{1}{2}\bar{H}_{CY})(\bar{z})},$ (56)
We also emphasize that these currents are cohomologies of the differential $D_{\vec{1},\vec{1}}$.
Taking the currents $J^{+}_{\omega}(\bar{z})$ and $J^{-}_{\dot{\omega}}$ and adding them to the currents of the algebra $SO(10)\times U(1)$, we extend the algebra $SO(10)\times U(1)$ to the algebra $E(6)$ and as a result obtain explicit expressions for all $78$ of the algebra $E(6)$, which is known to be considered as a possible Grand unified gauge group.
It should be noted that the emergence of the $E(6)$ symmetry in the right sector, firstly, ensures the self-consistency of the theory in connection with the presence of $N=1$ supersymmetry in the left sector. Secondly, it is both of these symmetries that are necessary for the Theory of Grand Unification.
Note that in the right sector, in addition to the vertex $V^{[10]}_{\vec{m}}=\exp{(i\vec{\Lambda}\cdot\vec{\Phi}+\vec{n}\cdot\vec{X}^{-})}$ which belong to the type $[V]$ we also find massless vertices $V^{[1]}_{\vec{m}}$ and $V^{[16]}_{\vec{m}}$ of the sixteen-dimensional $SO(10)$ representations.
These vertices are the $E(6)$ partners of $V^{[10]}_{\vec{m}}$, which also belong to the representation $27$, where
$V^{[1]}_{\vec{m}}=\exp{(-\bar{H}_{CY}+\vec{m}\cdot\vec{X}^{-})},$ (57)
and one more massless vertex that belongs to the $16$-dimensional spinor representation of $SO(10)$
$V^{[16]}_{\vec{m}}=\exp{(i\vec{\omega}\cdot\vec{\Phi}-\frac{1}{2}\bar{H}_{CY}+\vec{m}\cdot\vec{X}^{-})},$ (58)
These three sets of vertices are the cohomologies of the differentials $D_{\vec{1},\vec{1}}$ form a 27-dimensional representation of E(6).
The right massless sector also contains $\bar{27}$ representations of E(6), which are similarly constructed from dual representations of $SO(10)$ and whose vertices depend on $\vec{n}\in\Delta^{-}$ instead of $\vec{m}\in\Delta^{+}$ . The numbers of representations $27$ and $\bar{27}$ are equal to the number of points in the reflexive Batyrev polytopes for a given CY-manifold.